# Novel point cloud registration approach for noninvasive patient specific estimation of leaflet strain from 3D images of heart valves

### v1

@@ 4.2 Comparison with existing literature
Strain in TVs varies substantially across leaflets [19, 20]. In our study, patients with mild or greater regurgitation exhibited greater average areal strain and 1 principal strain in the anterior leaflet, but lower values in the posterior leaflet, compared with the trivial to moderate cohort. Across both normal and diseased groups, the posterior leaflet consistently experienced the greatest stretch, followed by the anterior and septal leaflets. This pattern aligns with findings reported by Spinner et al. [44]. However, Mathur et al. identified the anterior leaflet as having the highest areal strain [19]. Analysis at the individual leaflet level further showed that billowing leaflets tend to exhibit higher IQR in the 1 principal strain, suggesting greater strain heterogeneity. Strain analysis in larger cohorts is needed to verify these observations. Population-level PCA demonstrated that the normal valve, represented by the mean shape, has the lowest average strain and IQR. As valves deviate from this mean toward tethering or billowing phenotypes, both the average strain and the IQR increase.
In our MV studies, we observed higher average strain in the posterior leaflet compared with the anterior leaflet, consistent with the literature [15]. This trend was evident in our 1 principal strain calculations across pediatric and adult populations, and in both normal and diseased valves. In adults, El-Tallawi et al. [15] reported average isotropic strains of 0.076 and 0.093 in the anterior and posterior leaflets, respectively, for normal MVs (mean age 55.3 2.6 years). In our adult cohort (mean age 64.6 9.4 years), the average areal strain in the anterior and posterior leaflets was –0.215 vs. –0.107, and the average principal strain was 0.081 vs. 0.231. Our results indicate that the strain measured in the posterior leaflets is approximately 2.8 times higher than that in the anterior leaflets. This value is notably higher (two times) than that reported by El-Tallawi et al. We attribute this difference to the specific time points chosen for strain calculation. El-Tallawi et al. calculated strain from mid-systole to end-systole. In contrast, our study measured strain from mid-diastole to mid-systole. Since the leaflet experiences maximum pressure near mid-systole, calculating the strain across that wider time frame (mid-diastole to mid-systole) inherently captures a larger deformation range and, consequently, results in a higher calculated strain value.
@@ 4.3 Limitations and area for future work
5 Conclusions
We have developed a novel point-based registration approach for tracking leaflet motions across the diastole and systole frames. Our method provides a comparable strain estimation to that in FEA. Our approach was successfully applied to three groups of atrioventricular valves across pediatric and adult patients. The cohorts encompassed a broad spectrum of valve anatomy, from healthy mitral valves to those with restricted leaflets and pronounced billowing. Although we did not identify a unifying strain pattern that clearly distinguishes healthy from diseased valves, we observed a promising trend in which a higher median and IQR of strain distribution appears to be a strong marker of leaflet billowing. With further validation, this metric could serve as a novel quantitative tool to enhance prognostication of valvular disease.
Acknowledgments

### v2

@@ 4.2 Comparison with existing literature
Strain in TVs varies substantially across leaflets [25, 20]. In our study, patients with mild or greater regurgitation exhibited greater average areal strain and 1 principal strain in the anterior leaflet, but lower values in the posterior leaflet, compared with the trivial to moderate cohort. Across both normal and diseased groups, the posterior leaflet consistently experienced the greatest stretch, followed by the anterior and septal leaflets. This pattern aligns with findings reported by Spinner et al. [47]. However, Mathur et al. identified the anterior leaflet as having the highest areal strain [25]. Analysis at the individual leaflet level further showed that billowing leaflets tend to exhibit higher IQR in the 1 principal strain, suggesting greater strain heterogeneity. Strain analysis in larger cohorts is needed to verify these observations. Cohort-level PCA demonstrated that the normal valve, represented by the mean shape, has the lowest average strain and IQR. As valves deviate from this mean toward tethering or billowing phenotypes, both the average strain and the IQR increase.
In our MV studies, we observed higher average strain in the posterior leaflet compared with the anterior leaflet, consistent with the literature [10]. This trend was evident in our 1 principal strain calculations across pediatric and adult populations, and in both normal and diseased valves. In adults, El-Tallawi et al. [10] reported isotropic strain measurements in normal adult mitral valves and observed modest differences between anterior and posterior leaflets. In our adult cohort, posterior leaflet strain was 2.8 times higher than anterior leaflet strain, a more pronounced difference between leaflets than previously reported. We attribute this difference to variation in the cardiac phases selected for strain calculation. Prior work quantified strain over a narrower systolic interval, whereas our analysis spanned a broader interval from mid-diastole to mid-systole, which captures a larger deformation range and results in higher measured strain.
@@ 4.3 Limitations and area for future work
While we validate and demonstrate the ability of our method to accurately capture leaflet strains across diverse valve geometries, we were not powered to explore relationships between leaflet strain, geometry, and pathology; our findings in this domain are exploratory. Several additional technical considerations may influence the interpretation of leaflet strain, particularly the occurrence of negative areal strain (i.e., stretch less than 1) [38, 39, 29]. From a mechanical perspective, leaflet area would generally be expected to increase during systole as pressurization induces tensile loading and tissue stretch, although localized contraction or complex folding of the leaflet may occur in specific regions. When computing leaflet area of the segmental atrial surface between diastole and systole, however, we observed an overall negative percent change in area. This behavior could potentially reflect imaging and reconstruction limitations in addition to any true physiological effects. In systolic 3DE, coaptation zones and locally folded regions are not sufficiently resolved (Fig. 4- 6 panel A), which can preclude full capture of the native leaflet surface and lead to systematic underestimation of the true leaflet area. Because areal strain is derived directly from these surface measurements, such underestimation contributes to negative strain values. However, these effects arise from consistent constraints in image quality and segmentation and are therefore systematic across leaflets and patient populations, preserving the reproducibility of relative comparisons and regional trends. Although our pipeline is compatible with segmentations generated from multiple software platforms, differences in smoothing or surface regularization may introduce modest variability in local geometry and consequently in strain measurements. Continued improvements in image resolution and segmentation fidelity will be important for reducing these sources of bias and further enhancing quantitative accuracy of strain estimation.
@@
4.4 Concluding remarks
We have developed a novel point-based feature tracking approach for tracking leaflet motions across the diastole and systole frames. Our method provides a comparable strain estimation to that in FEA. Our approach was successfully applied to three groups of atrioventricular valves across pediatric and adult patients (a total of 31 cases). The cohorts encompassed a broad spectrum of valve anatomy, from healthy mitral valves to those with restricted leaflets and pronounced billowing. Although we did not identify a unifying strain pattern that clearly distinguishes healthy from diseased valves, we observed a promising trend in which a higher median and IQR of strain distribution appears to be a strong marker of leaflet billowing. With further validation, this metric could serve as a novel quantitative tool to enhance prognostication of valvular heart disease.
5 Acknowledgments
@@
CRediT authorship contribution statement
Wensi Wu: Conceptualization, Formal analysis, Funding acquisition, Investigation, Methodology, Visualization, Writing - original draft Matthew Daemer: Data curation, Writing – review & editing Jeffrey A. Weiss: Conceptualization, Funding acquisition, Writing – review & editing Alison M. Pouch: Conceptualization, Data curation, Funding acquisition, Writing – review & editing Matthew A. Jolley: Conceptualization, Funding acquisition, Writing – review & editing